### v2

# Information Loss and Cost in Symmetry Breaking

Abstract
We develop an algebraic and information–theoretic framework to characterize symmetry breaking of generalized, non-invertible symmetries in two spatial dimensions. The reduction of symmetry is modeled within the subfactor theory, where condensable Frobenius algebras play the role of subgroups in the categorical setting. This perspective naturally connects to the description of anyon condensation in topological phases of matter. Central to our approach are coarse-graining maps, or conditional expectations, which act as quantum channels projecting observables from a phase with higher symmetry onto one where the symmetry is partially or completely broken by condensation. By employing relative entropy as an entropic order parameter, we quantify the information loss induced by condensation and establish a universal bound governed by the Jones index, which is equal to the quantum dimension of the condensate. We illustrate the framework through explicit examples, including the toric code, abelian groups $\mathbb{Z}_{N}$, and the representation category $\mathrm{Rep}(S_{3})$, and show how dualities give rise to equivalence classes of condensation patterns. Our results forge new connections between operator algebras, tensor category theory, and quantum information in the study of generalized symmetries.
1 Introduction
Recent advances in the study of generalized global symmetries, including the so-called non-invertible symmetries, emphasize the role of extended topological operators that implement such symmetries. In quantum field theory (QFT), these generalized global symmetries are realized through topological defect operators whose composition rules are generally governed by fusion categories [1]. These symmetries impose stringent constraints on correlation functions and dualities and give rise to novel phases and topological orders [2, 3].
In QFT, a global symmetry usually involves an ordinary compact group acting on the algebra of fields. In such cases, patterns of symmetry breaking and the resulting phases are classified by subgroups of the original symmetry group. However, in theories exhibiting topological order and non abelian anyonic excitations, the relevant statistics are governed by the braid group, and the symmetries are more naturally described by the mathematical framework of fusion categories. These structures capture what is now referred to as categorical or non-invertible symmetries, extending the notion of symmetry beyond group actions. Unlike conventional group-like symmetries, non-invertible symmetries do not admit a notion of subgroups, which complicates the characterization of symmetry-broken phases. However, in these cases, algebraic objects known as condensable algebras can be thought of as a generalized analog of a subgroup, and the invariant subspaces under their action correspond to different gaugings/condensations or phases [4, 5].
Then, it is natural to investigate the patterns of symmetry breaking and the associated phases of systems with these generalized symmetries. There have been recent efforts to extend the Landau paradigm to non-invertible symmetries based on the idea of the SymTFT framework [6]. SymTFT characterizes the symmetry of a theory in ($d$ + 1) dimensions in terms of a ($d$ + 2)-dimensional topological quantum field theory. This allows us to classify gapped phases of the theory under consideration solely on the basis of the properties of the symmetry, including its anomalies. There has been progress for gapped phases in (1+1)-$d$ and (2+1)-$d$ [7, 8, 9, 10, 11, 12], and for gapless phases [13, 14, 15, 16, 17, 18].
Here, we propose an algebraic and information-theoretic framework to analyze symmetry-breaking patterns associated with generalized global symmetries. Although the algebraic approach to QFT focuses on local operator algebras [19, 20, 21, 22], the study of generalized global symmetries emphasizes global features and the role of extended operators. The way these topological objects implement the corresponding symmetry operations and their possible patterns of symmetry breaking are formulated in the language of tensor categories and, interestingly, also admit a natural and elegant formulation within subfactor theory of von Neumann algebras [23, 24, 25], that is, inclusions $\mathcal{N}\subset\mathcal{M}$, of von Neumann algebras, where $\mathcal{M}$ models a set of observables of a quantum system, while the subalgebra $\mathcal{N}$ describes a system with restricted observables or reduced symmetry.
Related ideas have been presented in [26, 27, 28], although our work provides a comprehensive treatment that combines concepts from condensed matter theory, theory of operator algebras, tensor category theory, and quantum information theory in the study of non-invertible symmetries. In condensed matter theory, we use the concept of anyon condensation [29]. In topologically ordered systems, anyon condensation refers to the process by which a subset of anyons—quasiparticle excitations in two-dimensional systems—condenses, giving rise to a new phase with modified topological order. Within the language of modular tensor categories, this phenomenon is captured by condensable algebras, a natural encoder of a symmetry sector in the case of non-invertible symmetries. These algebras identify the subset of bosonic anyons that can condense consistently with the ambient topological structure, thus encoding the symmetry-breaking patterns of the resulting phases [30].
Central to our proposal are the concepts of conditional expectations and coarse-grainings, which, in quantum information terms, are quantum channels, that is, completely positive maps that effectively ’project’ observables or states from $\mathcal{M}$ to $\mathcal{N}$ and vice versa.[31, 32]. In our setting, these channels encode how states and operators of the system under study resolve under different broken symmetry patterns of the full symmetry. Here we note that, while gauging and anyon condensation are often presented as distinct procedures, both are governed by the same algebraic structure, that is to say, a condensable algebra object. In operator-algebraic terms, this structure gives rise to a conditional expectation, which simultaneously implements gauging (projection onto invariant observables) and the symmetry reduction associated with condensation. Our proposal makes this dual perspective explicit.
In this framework, it is thus natural to ask which information-theoretic quantities can reliably quantify the loss of information incurred when transitioning from the original phase to the symmetry-broken phase with reduced symmetry. Inspired by the proposals in [34, 33], we introduce an entropic order parameter based on the quantum relative entropy to quantify the degree of symmetry breaking. Relative entropy plays a major role in quantum information theory in determining the information loss resulting from the processing of a quantum channel in a given state, that is, in distinguishing states. The definition of relative entropy in the context of von Neumann algebras was first introduced by Araki [35, 36]. A modern treatment and an exhaustive overview of the subsequent results can be found in [31, 32].
These entropic order parameters quantify the amount of distinguishability between arbitrary states in the theory and symmetry-averaged versions of these states for which the expectation values of the non-local operators implementing the symmetry are different. This reflects the information loss that occurs when observables are restricted to broken symmetry sectors. The symmetry-breaking patterns underlying these processes are defined in terms of condensable Frobenius algebras. Similar algebraic approaches as the one presented here have been used in other quantum information contexts such as in [37, 38]. Although many of our constructions admit a more rigorous formulation within the framework of von Neumann algebra theory, our focus is on highlighting the underlying physical principles. When necessary, we direct the reader to the appropriate mathematical literature for a more formal exposition.
The work is organized as follows. Section 2 introduces the algebraic preliminaries, reviewing the structure of von Neumann algebras, quantum states, and the definition of relative entropy, which serves as a key information-theoretic tool throughout the work. In Section 3 we present subfactor theory in the language of anyon condensation, interpreting symmetry-breaking transitions as inclusions of algebras $\mathcal{N}\subset\mathcal{M}$ and explaining the role of Frobenius algebras in encoding symmetry-broken sectors. Section 4 constructs the conditional expectation as a quantum channel that implements the restriction map from the uncondensed to the condensed phase and introduces the lifting map that reverses this process. In Section 5, the relative entropy between the original and lifted states is proposed as an entropic order parameter, and the framework is illustrated through detailed examples including the groups $\mathbb{Z}_{N}$, the toric code, and the representation category of $S_{3}$. We also include two appendices. In Appendix A we prove some properties of the conditional expectation of Section 3 and Appendix B discusses how the presence of dualities in the original theory implies dual condensation patterns leading to equivalent symmetry-broken patterns.
2 Background on von Neumann Algebras and states
To establish the basic terminology used throughout this work, we recall that quantum systems are described by their algebras of observables $\mathcal{M}$, which act on a Hilbert space $\mathcal{H}$. An operator algebra consists of a set of operators (or matrices) closed under linear combinations, products, and the *-operation (Hermitian conjugation in physics). The algebra also includes all scalar multiples of the identity operator. Formally, this reads as,
$\mathbf{1}\in{\mathcal{M}}\,,\hskip 14.22636pta,b\in{\mathcal{M}}\implies\hskip 14.22636pt\alpha a+\beta b\in{\cal M}\,,\hskip 14.22636ptab\in{\cal M}\,,\hskip 14.22636pta^{\dagger}\in{\cal M}\,$, (2.1)
with $\alpha,\beta\in\mathbb{C}$. The observables of the system are represented by self-adjoint elements $(a=a^{\dagger})$ in $\mathcal{M}$. We consider $\mathcal{M}\subset\mathcal{B}(\mathcal{H})$, where $\mathcal{B}(\mathcal{H})$ denotes the algebra of bounded operators in $\mathcal{H}$. $\mathcal{M}$ is called a von Neumann algebra if it satisfies the condition,
$\mathcal{M}=\mathcal{M}^{\prime\prime}\,$ (2.2)
where the prime refers to the commutant,
$\mathcal{M}^{\prime}=\{b/\,\,[b,a]=0,\,\forall a\in\mathcal{M}\}.$ (2.3)
In finite dimensions, von Neumann algebras can be expressed in terms of matrix algebras
${\mathcal{M}}=\bigoplus_{i}\,M^{(i)}_{m_{i}}\otimes 1_{n_{i}\times n_{i}}\,,\quad{\mathcal{M}}^{\prime}=\bigoplus_{i}\,1_{m_{i}\times m_{i}}\otimes N^{(i)}_{n_{i}}\,$, (2.4)
where $M^{(i)}_{m_{i}}$ and $N^{(i)}_{n_{i}}$ are $m_{i}\times m_{i}$ and $n_{i}\times n_{i}$ arbitrary complex matrices, respectively. The standard representation of the algebra $\mathcal{M}$} is the one obtained by choosing all the $N^{(i)}_{n_{i}}=1$. This amounts to the fact that the minimal projectors of the algebra (projectors such that there are no other projectors into smaller subspaces) are of dimension $1$.
A von Neumann algebra $\mathcal{M}$ is called a factor if $\mathcal{M}\cap\mathcal{M}^{\prime}=\mathbb{C}\cdot\mathbf{1}$, i.e. its center is trivial. Every von Neumann algebra can be decomposed as a direct sum (or, more generally, a direct integral) of factors. A subfactor is a unital inclusion $\mathcal{N}\subset\mathcal{M}$ of von Neumann algebras where both $\mathcal{N}$ and $\mathcal{M}$ are factors (i.e., they have trivial centers). A subfactor is irreducible if $\mathcal{N}^{\prime}\cap\mathcal{M}=\mathbb{C}\,\cdot\mathbf{1}$.
Von Neumann factors are classified into three types, denoted I, II, and III. In this work, we focus on Type I factors. These are von Neumann algebras that, for a given Hilbert space $\mathcal{H}$, are isomorphic to the algebra of bounded operators $\mathcal{B}(\mathcal{H})$. The classification within Type I is based on the dimension of $\mathcal{H}$: for example, a Type In factor is isomorphic to the matrix algebra $M_{n}(\mathbb{C})$. More generally, every finite-dimensional von Neumann algebra decomposes as a direct sum of Type I factors.
In this framework, quantum states are described by expectation value functionals, which assign to each operator $a\in\mathcal{M}$ a complex number representing its expectation value. Formally, a state is a linear map $\omega:{\mathcal{M}}\rightarrow{\mathbb{C}}$. These functionals are positive, meaning that $\omega(a)\geq 0$ for every self-adjoint operator $a\in\mathcal{M}$ with non-negative spectrum, and normalized, so that the total probability assigned to a collection of mutually orthogonal projectors summing to the identity is one. Then a state is defined as
$\omega(\alpha a+\beta b)=\alpha\,\omega(a)+\beta\,\omega(b)\,,\hskip 14.22636pt\omega(aa^{\dagger})\geq 0\,,\hskip 14.22636pt\omega(\mathbf{1})=1\,.$ (2.5)
Such states are referred to as normal states on $\mathcal{M}$, and describe the physical states of the quantum system.
In finite-dimensional algebras $\mathcal{M}$, each normal state $\omega$ can be uniquely represented by a density matrix $\rho_{\omega}\in{\mathcal{M}}$ given by the relation
$\omega(a)={\rm tr}\rho_{\omega}a\,,\,\,\,\forall a\in{\mathcal{M}}\,$, (2.6)
where the trace is taken in the standard representation of the algebra. In this representation, $\rho_{\omega}$ takes the block diagonal form
$\rho_{\omega}=\left(\begin{array}[]{ccc}p^{\omega}_{1}\,\rho^{(1)}&0&0\\ 0&p^{\omega}_{2}\,\rho^{(2)}&0\\ 0&0&\ddots\end{array}\right)\,$, (2.7)
where the coefficients $p^{\omega}_{i}\geq 0$, with $\sum_{i}p^{\omega}_{i}=1$, represent probabilities, and each $\rho^{(i)}$ is a $m_{i}\times m_{i}$ non negative hermitian matrix with trace one.
Interestingly, in the finite-dimensional case, a state $\omega\in\mathcal{M}$ corresponds to a specific operator within $\mathcal{M}$ itself, namely the density matrix $\rho_{\omega}$. In contrast, for infinite-dimensional algebras, the state is no longer represented by a density matrix but rather by a more general object that fulfills the same role of generating well-defined expectation values. This object is known as the modular operator.
The relative entropy is a central quantity in quantum information theory. For matrix algebras, and more generally type I algebras, it is defined by
$S_{\mathcal{M}}(\rho\,|\,\sigma)={\rm tr}_{\mathcal{M}}[\rho(\log\rho-\log\sigma)]\,$, (2.8)
where $\textrm{tr}_{\mathcal{M}}$ denotes the canonical trace associated with the algebra $\mathcal{M}$, and $\rho$, $\sigma$ are the two density matrices associated with the two states in comparison; see [31]. Relative entropy provides a useful measure of how distinguishable two quantum states are. Given a state $\sigma$, the probability $p$ that the outcomes of $N$ independent measurements are consistent with the predictions of a different state $\rho$ decays exponentially with $N$, following the asymptotic behavior $p\sim e^{-NS(\rho|\sigma)}$.
3 Subfactor Theory, Anyon Condensation and Entropic Order parameters
Let us return to the discussion in Section 2 regarding a subfactor, which, recall, refers to a unital inclusion $\mathcal{N}\subset\mathcal{M}$ of von Neumann algebras, where both $\mathcal{N}$ and $\mathcal{M}$ are factors.
In the physical context, we are considering a quantum mechanical system that is described by an algebra $\mathcal{M}$ of operators, and the dynamical variables (or observables) correspond to the self-adjoint elements. We also assume that the level of observation does not allow to access the expectation value of all observables, but only a "small" part of them represented by the subalgebra $\mathcal{N}$ of the full algebra $\mathcal{M}$.
A fundamental notion in subfactor theory is that of a conditional expectation $\varepsilon:\mathcal{M}\to\mathcal{N}$. It is a unital, completely positive map that acts as a projection from $\mathcal{M}$ onto $\mathcal{N}$, and satisfies the bimodule property:
$\varepsilon(n_{1}mn_{2})=n_{1}\varepsilon(m)n_{2}\,\quad\forall n_{1},n_{2}\in\mathcal{N},m\in\mathcal{M}.$
In general, a given inclusion $\mathcal{N}\subset\mathcal{M}$ may not admit any conditional expectation at all. However, in the case of an irreducible inclusion, if a normal conditional expectation exists, then it is unique. A key result states that for systems with finite-dimensional Hilbert space $\mathcal{H}$, a conditional expectation $\varepsilon$ can be expressed as:
$\varepsilon(m)=\sum_{i=1}^{N}\mathbb{K}_{i}\,m\,\mathbb{K}^{\dagger}_{i}$ (3.1)
where $m\in\mathcal{M}$, and the $\mathbb{K}_{i}$ are the so-called Kraus operators. In this form, a conditional expectation appears as a specific example of a quantum channel. Complementary to this is the concept of coarse-graining. This is a positive linear embedding map or quantum channel $\alpha:\mathcal{N}\to\mathcal{M}$ that also provides partial information of the total quantum system $\mathcal{M}$.
Finally, it is necessary to introduce the notion of the index $\lambda$ of a subfactor $\mathcal{N}\subset\mathcal{M}$, a quantity that links inclusions of von Neumann algebras to the concrete realizations of quantum channels in quantum information theory. For an irreducible subfactor $\mathcal{N}\subset\mathcal{M}$, and assuming the existence of a unique conditional expectation $\varepsilon:\mathcal{M}\to\mathcal{N}$, the Jones (or Jones–Kosaki–Longo) index is defined as [39, 40, 41, 42]
$\lambda=[\mathcal{M}:\mathcal{N}]:=[\sup\{\lambda>0:\varepsilon(m)\geq\lambda\,m\,\textrm{ for all }m\in\mathcal{M}_{+}\}]^{-1}.$ (3.2)
This definition makes use of the Pimsner–Popa inequality [43], where $\mathcal{M}_{+}$ denotes the positive cone of the von Neumann algebra $\mathcal{M}$. Explicitly,
$\mathcal{M}_{+}:=\{m\in\mathcal{M}|m=m^{*},{\rm and}\;{\rm spec}(m)\subseteq[0,\infty)\}\,$, (3.3)
that is, $\mathcal{M}_{+}$ consists of all positive semidefinite operators in $\mathcal{M}$. If no such a conditional expectation exists, the index is defined to be $\mathcal{M}:\mathcal{N}]=\infty$. The index $\lambda=[\mathcal{M}:\mathcal{N}]$ quantifies the relative "size" of $\mathcal{M}$ compared to $\mathcal{N}$, analogous to the index of a subgroup but enriched by the analytic structure of operator algebras. It also provides an upper bound on the dimension of $\mathcal{M}$ when viewed as a module over $\mathcal{N}$
Finally, it is worth mentioning that, motivated by classical ergodic theory, Connes and Störmer [44] introduced the relative entropy $H(\mathcal{M}|\mathcal{N})$ for inclusion $\mathcal{N}\subset\mathcal{M}$ of finite (dimensional) algebras. The connection between these two quantities was established by Pimsner and Popa [43] who proved the general relation
$H(\mathcal{M}|\mathcal{N})\leq\log[\mathcal{M}:\mathcal{N}]\,.$ (3.4)
In this work, we are interested in the patterns of symmetry breaking in theories with non-invertible symmetries. Subfactor theory offers a rich mathematical setting to describe these phenomena, as subfactors serve as universal hosts for generalized quantum symmetries [45]. Here, we explicitly address this idea using the language of topological symmetry breaking and anyon condensation, whose main mathematical description is given in terms of condensable Frobenius algebras [8].
Concretely, let us consider a system of anyons with a generalized noninvertible symmetry governed by a Hamiltonian such that, by tuning one of its parameters, an anyon with charge $a$ undergoes condensation, that is, a new vacuum is formed, in which the topological charge $a$ is indistinguishable (this requires $a$ to be a boson). The resulting condensate breaks the original symmetry encoded in the topological excitation spectrum of the system. Before condensation, the superselection sectors, that is, the anyon types, are described by an algebra of operators $\mathcal{A}$ corresponding to the anyon species of a modular tensor category describing the uncondensed phase, with fusion rules determined by the decomposition of the tensor products of these representations. The condensation process effectively reduces $\mathcal{A}$ to a smaller subalgebra $\mathcal{T}$, whose excitations are described by operators given by irreducible representations of a fusion category describing anyonic excitations after condensation. In this transition, certain operators of $\mathcal{A}$ become decomposable in terms of elements of $\mathcal{T}$. Furthermore, some operators that were nonequivalent in $\mathcal{A}$ may become equivalent in $\mathcal{T}$, and thus are identified in the broken phase.
The condensation process and the subsequent reduction in the amount of symmetry of the system are described in terms of an algebraic object known as Frobenius or condensable algebra $\mathbb{A}$. The condensable algebra is the sum of simple objects of the uncondensed category with associated algebra $\mathcal{A}$. From an algebraic operator perspective, a condensable algebra object can be thought of as a generalized analog of a subgroup, and the invariant subspace under its action corresponds to different gaugings/condensations or phases [25]. As will be shown, $\mathbb{A}$ defines a conditional expectation that acts as a projector from the pre-condensed algebra to the post-condensed smaller subalgebra $\varepsilon_{\mathbb{A}}:\mathcal{A}\to\mathcal{T}$, which amounts to the inclusion of subfactors $\mathcal{T}\subset\mathcal{A}$. That is, we decompose the larger algebra $\mathcal{A}$ into elements of $\mathcal{T}$. In the formalism of anyon condensation developed in  [46, 47] that we follow from here on, a condensation algebra gauging or projection is implemented in terms of the so-called restriction map, which imposes a set of conditions that relate the particle excitations in $\mathcal{A}$ to those in the condensed phase $\mathcal{T}$. In general, the operators in $\mathcal{T}$ refer to elements of a fusion category rather than a modular one, since it may include excitations that are confined by the surrounding condensate. By projecting these confined excitations, one obtains the deconfined sector of condensed theory, described by an algebra of operators $\mathcal{U}$ with the fusion rules determined now as a modular tensor category.
In this work, since we are interested in the patterns of symmetry breaking in theories with generalized symmetries, we focus solely on determining the structure of $\mathcal{T}$, given the original algebra $\mathcal{A}$ and basic information about the condensate, such as the identification of the bosonic particles that condense. In this sense, we do not address the full topological phase transition from $\mathcal{A}$ to $\mathcal{U}$, but instead concentrate solely on the symmetry breaking from $\mathcal{A}$ to $\mathcal{T}$.  [^1] We note that the mathematical literature regarding anyon condensation uses category theory and the Frobenius algebra formalism and $Q$-systems [45]. A $Q$-system is a mathematical object that encodes a $C^{*}$-Frobenius algebra in a fusion category consisting of a triple $(\mathbb{A},\mu,\Delta)$ where: $\mathbb{A}$ is an object in a fusion category (for example, an MTC), $\mu:\mathbb{A}\otimes\mathbb{A}\to\mathbb{A}$ is a multiplication morphism and $\Delta:\mathbb{A}\to\mathbb{A}\otimes\mathbb{A}$ is a co-multiplication morphism. These morphisms satisfy associativity, co-associativity, and Frobenius conditions ensuring $\mathbb{A}$ forms a Frobenius algebra. In the theory of anyon condensation, $\mathbb{A}$ represents the condensable algebra. Therefore, if an algebra $\mathcal{A}$ corresponding to MTC describes the uncondensed phase, the condensation process leads to a subalgebra $\mathcal{T}$, induced by the $Q$-system..
Therefore, following [46, 47, 48] we determine the algebra $\mathcal{T}$, characterized by particle labels $\phi,r,s,t,..$, along with a set of branching (or restriction) coefficients $n_{a}^{t}$. These coefficients describe how the excitations of the original uncondensed algebra $\mathcal{A}$ decompose when certain bosonic sectors $\gamma_{i}$ condense to form a new vacuum sector $\phi$ in $\mathcal{T}$:
$\phi=1+\gamma_{1}+\cdots+\gamma_{n}\,.$ (3.5)
The branching coefficients implement the so-called restriction map, which characterizes the new phase:
$a\to\sum_{t\in\mathcal{T}}n^{t}_{a}\,t\,\qquad\forall a\in\mathcal{A}\,\qquad\text{(restriction)}$ (3.6)
where $n_{a}^{t}\in\mathbb{Z}_{\geq 0}$. It is assumed that $n^{t}_{a}=n^{\bar{t}}_{\bar{a}}$, where bars denote antiparticles. If more than one particle $t$ appears on the right-hand side of (3.6), we say that “$a$ splits into $\sum_{t}n_{a}^{t}t$”. The condensed particles satisfy the additional condition that $n^{\phi}_{\gamma_{i}}\neq 0$, indicating that the $\gamma_{i}$ sector contributes to the vacuum $\phi$ in the condensed phase $\mathcal{T}$.
In this setting, one can interpret $a$ as an operator of the algebra $\mathcal{A}$, which may decompose into smaller modules when the action is restricted to a subalgebra $\mathcal{T}\subset\mathcal{A}$. The same branching coefficients also define the adjoint of the restriction map, commonly referred to as the lift map
$t\to\sum_{a\in\mathcal{A}}n^{t}_{a}\,a\,\qquad\forall t\in\mathcal{T}\,\qquad\text{(lift)}$ (3.7)
Consistency conditions on the branching coefficients [48] impose two constraints on the quantum dimensions of particle excitations defined by algebras $\mathcal{A}$ and $\mathcal{T}$
$d_{a}$ $=$ $\sum_{t\in\mathcal{T}}n_{a}^{t}d_{t},\quad\forall~a\in\mathcal{A},$ (3.8a)
$d_{t}$ $=$ $\frac{1}{q}\sum_{a\in\mathcal{A}}n^{t}_{a}d_{a},\quad\forall~t\in\mathcal{T},$ (3.8b)
where $q$ denotes the quantum dimension of the condensed vacuum, given by
$q:=\sum_{a\in\mathcal{A}}n^{\phi}_{a}d_{a}=D_{\mathcal{A}}^{2}/D_{\mathcal{T}}^{2}\,$, (3.9)
where
$D^{2}_{\mathcal{A}}=\sum_{a\in\mathcal{A}}\,d_{a}^{2}\,,\quad D^{2}_{\mathcal{T}}=\sum_{t\in\mathcal{T}}\,d_{t}^{2}\,$, (3.10)
are the total (Perron-Frobenius) quantum dimensions of $\mathcal{A}$ and $\mathcal{T}$ respectively. Subfactors offer a natural operator-algebraic framework for describing anyon condensation: in this language, the Jones index $\lambda=[\mathcal{A}:\mathcal{T}]$ corresponds to the quantum dimension of the condensate $q$.
Therefore, it is expected that the above formalism can be naturally expressed in terms of a conditional expectation $\varepsilon:\mathcal{A}\to\mathcal{T}$ implementing the restriction map (3.6). This map acts as a projection, resolving elements of $\mathcal{A}$ in terms of operators in the subalgebra $\mathcal{T}$.
We can summarize the previous discussion as [25]:
$\text{Subfactor $\mathcal{T}\subset\mathcal{A}$}\longleftrightarrow\text{Pattern of symmetry breaking $\equiv\mathbb{A}$}$
It is worth establishing a connection between the framework discussed above and the gauging of non-invertible symmetries. In fact, the existence of a conditional expectation associated with a condensable algebra $\mathbb{A}$ is closely connected to the gauging of a non-invertible symmetry. This correspondence has been explored in detail in [51, 52, 4, 5]. In this context, the gauging of the symmetry is defined via an algebra object $\mathbb{A}=\sum_{a\in\mathcal{A}}{\rm n}_{a}\,a$, which is a weighted sum over simple objects $a\in\mathcal{A}$. Since $a$ corresponds to simple topological operators, the gauging of a partition function can be understood as summing over all configurations of topological labeled in $\mathbb{A}$. In our language, $\mathbb{A}$ is defined in terms of the vacuum $\phi$ and the coefficients ${\rm n}_{a}\equiv n_{a}^{\phi}$ that determine how each anyon $a$ contributes to the condensate.
3.1 Conditional expectation for the restriction and lifting maps
In this work, we are interested in analyzing different patterns of broken symmetry in systems with generalized symmetries. This amounts to considering different gaugings defined through condensable algebra objects $\mathbb{A}$. Here, from a quantum information perspective, we describe the gauging of a generalized symmetry and its inversion using the language of anyon condensation. While gauging and anyon condensation are distinct procedures, both are governed by the same algebraic structure: a condensable algebra object. In operator-algebraic terms, this structure gives rise to a conditional expectation, which simultaneously implements a gauging map (projection onto invariant observables) and the condensation map encoding the symmetry reduction associated with condensation.
Our framework is thus based on this unifying perspective, making use of the restriction and lifting maps to build the quantum channels (conditional expectations) yielding density operators corresponding to different gaugings-a.k.a. broken symmetry patterns-as
${\rm Condensable\,Algebra\,\mathbb{A}}\to n^{t}_{a}\implies\varepsilon_{\mathbb{A}}:\rho\in\mathcal{A}\longrightarrow\rho_{\mathbb{A}}\in\mathcal{T}$
The density operator $\rho_{\mathbb{A}}=\varepsilon_{\mathbb{A}}(\rho)$ represents a gauged/averaged version of the density operator $\rho$ w.r.t. the subfactor structure defined by the Frobenius algebra $\mathbb{A}$. The coarse-graining $\alpha$ that implements the lifting map, expresses the gauged density operator in terms of the original algebra $\mathcal{A}$
$\alpha:\rho_{\mathbb{A}}\in\mathcal{T}\longrightarrow\tilde{\rho}\in\mathcal{A}\,.$
In this sense, we have discussed above how a subfactor inclusion $\mathcal{N}\subset\mathcal{M}$ defined by a conditional expectation $\varepsilon:\mathcal{M}\rightarrow\mathcal{N}$ can be implemented via Kraus operators, as shown in Eq. (3.1). Turning to the subfactor inclusion $\mathcal{T}\subset\mathcal{A}$, here we consider how to build the corresponding conditional expectation in terms of Kraus operators $\mathbb{K}_{a}$ as,
$\varepsilon_{\mathbb{A}}(\rho)=\sum_{a\in\mathcal{A}}\mathbb{K}_{a}\,\rho\,\mathbb{K}^{\dagger}_{a}\,$, (3.11)
in order to implement the gauging associated with a condensation algebra $\mathbb{A}$  [^2] In this sense, one might interpret $\varepsilon$ as the condensation defect operator [53]. A condensation defect can be understood as the insertion of a projection operator within a decomposing QFT, particularly when the defect is associated with the higher gauging of a global symmetry. Thus, the condensation defect, analogously to $\varepsilon$, projects states in the theory onto specific symmetry sectors..
First, we note that the structure of the state space (an operator space) of an anyonic system prohibits superpositions between different topological charges due to anyonic superselection rules [54, 55]. These superselection rules require that an anyonic state be represented by a density operator that is a direct sum of sectors $\bigoplus_{a}\rho_{a}$. As a result, we consider an Abelian algebra $\mathcal{A}$ representing the uncondensed phase. For this algebra, the fusion rules of the non-local symmetry operators can be diagonalized in terms of a set of projectors $\Pi_{a}$ on the states of $\mathcal{A}$ with charge $a$ in such a way that in the standard representation of the algebra, a general state is given by the density matrix
$\rho=\sum_{a\in\mathcal{A}}\,p_{a}\,\Pi_{a}\,$, (3.12)
where projectors $\Pi_{a}$ satisfy the following conditions
$\Pi_{a}^{\dagger}=\Pi_{a},\quad\Pi_{a}\Pi_{b}=\delta_{ab}\Pi_{a}$ (3.13)
and $\{p_{a}\}$ are the corresponding probabilities $p_{a}\equiv{\rm tr}(\rho\,\Pi_{a})$ such that $\sum_{a}p_{a}=1$.  [^3] As an example, at infinite temperature $\beta\to 0$ one has that [50] $p_{a}=\frac{N_{a}\,d_{a}}{\sum_{b\in\mathcal{A}}N_{b}\,d_{b}}\,$, (3.14) with $N_{a}$ referring to the multiplicity of the irrep $a$ and $d_{a}$ the corresponding quantum dimensions.
To construct the conditional expectation $\varepsilon_{\mathbb{A}}:\mathcal{A}\to\mathcal{T}$ that implements the restriction map associated with a condensable algebra $\mathbb{A}$, we define the corresponding Kraus operators using the branching coefficients $n^{t}_{a}$ as follows,
$\mathbb{K}_{a}$ $=\sum_{t\in\mathcal{T}}\,\sqrt{\,n^{t}_{a}\left(\frac{d_{t}}{d_{a}}\right)}\,\Lambda_{t,a}\,,\quad a\in\mathcal{A},\,\quad t\in\mathcal{T}\,$, (3.15)
where $\Lambda_{t,a}$, and its adjoint $\Lambda_{a,t}=\Lambda_{t,a}^{\dagger}$ are operators that satisfy
$\Lambda_{t,a}\Pi_{b}=\delta_{ab}\Lambda_{ta},\quad\Lambda_{t,a}\Lambda_{b,s}=\delta_{ab}\delta_{ts}\Pi_{t}\,.$ (3.16)
With this definition, the action of $\varepsilon_{\mathbb{A}}$ on $\rho$ is given by
$\varepsilon_{\mathbb{A}}(\rho)$ $=\sum_{a\in\mathcal{A}}\,\mathbb{K}_{a}\,\rho\,\mathbb{K}_{a}^{\dagger}=\sum_{a\in\mathcal{A}}\,\mathbb{K}_{a}\,\left(\sum_{a\in\mathcal{A}}\,p_{a}\,\Pi_{A}\right)\,\mathbb{K}_{a}^{\dagger}$
$=\sum_{t,a}\,n^{t}_{a}\left(\frac{d_{t}}{d_{a}}\right)\,p_{a}\,\Pi_{t}=\sum_{t\in\mathcal{T}}\,p_{t}\,\Pi_{t}\,$ (3.17)
where the probabilities in the condensed phase are
$p_{t}=\sum_{a\in\mathcal{A}}\,n^{t}_{a}\left(\frac{d_{t}}{d_{a}}\right)\,p_{a}\,.$ (3.18)
So defined, the conditional expectation $\varepsilon_{\mathbb{A}}$ is trace preserving, idempotent (so acts as a projector) and fulfills the bimodule property. The proofs of the latter are presented in Appendix A. The trace-preserving property of the conditional expectation $\varepsilon_{\mathbb{A}}$ can be easily verified as follows:
${\rm tr}\,\varepsilon_{\mathbb{A}}(\rho)=\sum_{t\in\mathcal{T}}\,p_{t}=\sum_{a\in\mathcal{A}}\frac{1}{d_{a}}\,p_{a}\,\sum_{t\in\mathcal{T}}\,n^{t}_{a}\,d_{t}=\sum_{a\in\mathcal{A}}\,p_{a}={\rm tr}\,\rho\,$, (3.19)
where in the final step we used the quantum dimension identity (3.8a),
$d_{a}=\sum_{t\in\mathcal{T}}\,n_{a}^{t}\,d_{t}\,.$ (3.20)
Furthermore, the Kraus operator condition is satisfied:
$\sum_{a\in\mathcal{A}}\,\mathbb{K}_{a}^{\dagger}\,\mathbb{K}_{a}$ $=\sum_{a\in\mathcal{A}}\sum_{t,s\in\mathcal{T}}\,\sqrt{n^{t}_{a}\,n^{s}_{a}\frac{d_{t}d_{s}}{d_{a}^{2}}}\,\Lambda_{a,t}\Lambda_{s,a}=\sum_{a\in\mathcal{A}}\,\frac{1}{d_{a}}\,\left(\sum_{t\in\mathcal{T}}n^{t}_{a}\,d_{t}\right)\,\Pi_{a}$ (3.21)
$=\sum_{a\in\mathcal{A}}\,\Pi_{a}=\mathbb{I}_{\mathcal{A}}\,.$
In addition to this, one can define the coarse-graining quantum channel $\alpha:\mathcal{T}\to\mathcal{A}$ that implements the lifting map through:
$\mathbb{L}_{a}=\sum_{t\in\mathcal{T}}\,\sqrt{\frac{n^{t}_{a}}{\lambda}\left(\frac{d_{a}}{d_{t}}\right)}\,\Lambda_{a,t}\,\quad a\in\mathcal{A}\,\quad t\in\mathcal{T}\,$, (3.22)
where
$\lambda=[\mathcal{A}:\mathcal{T}]=\sum_{a\in\mathcal{A}}n^{\phi}_{a}d_{a}=q\,$, (3.23)
is the Jones-index of the subfactor inclusion. With this, the lifted density matrix $\tilde{\rho}$ is given by
$\widetilde{\rho}$ $=\alpha(\rho_{\mathbb{A}})=\sum_{a\in\mathcal{A}}\,\mathbb{L}_{a}\,\varepsilon_{\mathbb{A}}(\rho)\,\mathbb{L}_{a}^{\dagger}=\sum_{a\in\mathcal{A}}\,\mathbb{L}_{a}\,\left(\sum_{t\in\mathcal{T}}\,p_{t}\,\Pi_{t}\right)\,\mathbb{L}_{a}^{\dagger}$ (3.24)
$=\frac{1}{\lambda}\sum_{t,a}\,n^{t}_{a}\left(\frac{d_{a}}{d_{t}}\right)\,p_{t}\,\Pi_{a}=\sum_{a\in\mathcal{A}}\,\widetilde{p}_{a}\,\Pi_{a}\,$,
where the lifted probabilities $\tilde{p}_{a}$ are given by
$\widetilde{p}_{a}$ $=\frac{1}{\lambda}\sum_{t\in\mathcal{T}}\,n^{t}_{a}\left(\frac{d_{a}}{d_{t}}\right)\,p_{t}\,.$ (3.25)
Although the operators $\mathbb{L}_{a}$ do not form a set of Kraus operators, they nonetheless preserve the trace of $\tilde{\rho}$, that is, $\sum_{a\in\mathcal{A}}\tilde{p}_{a}=1$, as ensured by the identity (3.8b). Moreover, the operators $\mathbb{L}_{a}$ satisfy an interesting trace identity:
$\sum_{a\in\mathcal{A}}{\rm tr}\left(\mathbb{L}_{a}\mathbb{L}_{a}^{\dagger}\right)=|\mathcal{T}|\,$, (3.26)
where $|\mathcal{T}|$ denotes the number of particle types (or simple objects) in the condensed theory.
As a remark, we notice that Equation (3.25) can also be written as
$\widetilde{p}_{a}$ $=\frac{1}{\lambda}\sum_{b\in\mathcal{A}}\,M^{\prime}_{ab}\left(\frac{d_{a}}{d_{b}}\right)\,p_{b}\,$, (3.27)
with the matrix
$M^{\prime}_{ab}=\sum_{t\in\mathcal{T}}n^{t}_{a}n^{t}_{b}$ (3.28)
defined in [48]. Restricting the summation in this equation to the unconfined particles $u\in\mathcal{U}$, one obtains the matrix $M_{ab}=\sum_{t\in\mathcal{U}}n^{t}_{a}\,n^{t}_{b}$ that plays a central role in the construction of the MTC $\mathcal{U}$ [48] that, as explained above, is not the purpose of our work.
From the perspective of operator algebras, our lifting map $\alpha:\,\mathcal{T}\to\mathcal{A}\,$ can be viewed as a state-level analogue of the well-known $\alpha$-induction procedure [56, 57], which lifts superselection sectors using the same branching data that enter our formulas; here we obtain its decategorified-braiding-blind counterpart, acting instead on states.
3.2 Relative Entropy as order parameter
With the previous constructions in place, we are now in a position to entropically quantify the information loss due to symmetry breaking after the condensation process using the relative entropy defined in (2.8). Although the computation of this quantity in QFT is not straightforward, in [34] the authors noted that this simplifies when, as in our case, $\mathcal{A}$ and $\mathcal{T}$ are small algebras for which the expectation values of the nonlocal operators are known. That is, the expectation values of the projectors $\Pi_{a}$ in both states $\rho$ and $\tilde{\rho}$ are
$p_{a}\equiv{\rm tr}(\rho\,\Pi_{a})\,,\qquad\tilde{p}_{a}\equiv{\rm tr}(\tilde{\rho}\,\Pi_{a})\;,$ (3.29)
one finds that the relative entropy is given by
$S_{\mathcal{A}}(\rho\,|\,\tilde{\rho})={\rm tr}_{\mathcal{A}}[\rho(\log\rho-\log\tilde{\rho}]\,=\sum\limits_{a\in\mathcal{A}}\,p_{a}\,\log\frac{p_{a}}{\tilde{p}_{a}}.$ (3.30)
This quantity provides a way to compare a density matrix in the algebra $\mathcal{A}$ before and after the condensation process. In this sense, it serves as an entropic order parameter for condensation, in line with the framework proposed in Ref. [34].
An equivalent expression, derived using Eq. (3.27), is
$S_{\mathcal{A}}(\rho|\tilde{\rho})=\log\lambda-H(p)-\sum\limits_{a\in\mathcal{A}}\,p_{a}\,\log\left[\sum\limits_{b\in\mathcal{A}}\,M^{\prime}_{ab}\left(\frac{d_{a}}{d_{b}}\right)\,p_{b}\right]\;.$ (3.31)
where $H(p)$ is the Shannon entropy of the initial probability distribution $\{p_{a}\}$, defined as
$H(p)=-\sum_{a\in\mathcal{A}}p_{a}\log p_{a}\,.$ (3.32)
The matrix $M^{\prime}_{ab}$ in our coarse-graining formula (3.31) coincide with the coupling matrices that appear in the $\alpha$-induction framework [56, 57]. There, such matrices encode modular invariants governing the extension of superselection sectors. In our setting, the same data control the probabilities $\tilde{p}_{a}$ of the lifted states and hence the loss of distinguishability as measured by relative entropy. Namely, for well-defined probability distributions and restriction-lifting maps, there is a bound
$S_{\mathcal{A}}(\rho|\,\tilde{\rho})\leq\log\lambda$ (3.33)
for the maximum amount of information loss associated to a possible gauging or symmetry broken phase given in terms of the Jones index of the inclusion.
4 Examples
[…]
4.1 Abelian group $\mathbb{Z}_{N}$.
Let us first focus on the group $\mathbb{Z}_{N}$, for which we consider the condensable algebra,
$\mathbb{A}=\sum\limits_{r}\,r\,$ (4.1)
with $r\in{\rm irreps}\,G$. In this case, $\mathcal{T}=\{\phi\}$ with $n^{\phi}_{r}=1,\,\forall\,r$ and $p_{\phi}=1$. The index $\lambda$ is given by
$\lambda=|G|\,.$ (4.2)
In the mathematical literature, the condensable algebra $\mathbb{A}$ above is known as a Lagrangian algebra. In terms of branching coefficients, a Lagrangian algebra is defined through coefficients that satisfy i) $n^{\phi}_{a}\neq 0,\,\forall\,a\in\mathcal{A}$, and ii) that the quantum dimension of the condensed vacuum $q$, that is, the index $\lambda$ is
$\lambda=\sum_{a\in\mathcal{A}}\,n^{\phi}_{a}\,d_{a}=\mathcal{D}_{\mathcal{A}}^{2}\,\implies\mathcal{D}_{\mathcal{T}}^{2}=1\,.$ (4.3)
Thus, the gauging induced by a Lagrangian algebra $\mathbb{A}$ and the related pattern of symmetry breaking result in a new vacuum $\phi$ in which all the non-trivial anyonic charges $a$ have condensed. We will see another example of Lagrangian algebra in the case of non-invertible symmetries.
Said this, in for the case of the group $\mathbb{Z}_{N}$, defining
${\rm n}^{\phi}=\begin{pmatrix}n_{1}^{\phi}&n_{2}^{\phi}&\cdots&n_{N}^{\phi}\end{pmatrix}=\begin{pmatrix}1&1&\cdots&1\end{pmatrix}$ (4.4)
we note that
$M^{\prime}=({\rm n}^{\phi})^{\rm T}{{\rm n}^{\phi}}=\begin{pmatrix}1&\cdots&1\\ \vdots&\ddots&\vdots\\ 1&\cdots&1\\ \end{pmatrix}$ (4.5)
From this matrix $M^{\prime}_{rs}$ it is straightforward to see that
$\sum_{s}\,M^{\prime}_{rs}\,p_{s}=p_{\phi}=1\,\quad\forall r\,$, (4.6)
which implies that
$S_{\mathcal{A}}(\rho|\,\tilde{\rho})=\log\lambda+\sum\limits_{r}\,p_{r}\,\log p_{r}=\log|G|-H(p)\,$, (4.7)
a result that was first obtained in [34]. It is worth noting that the lifted density matrix $\tilde{\rho}$ is proportional to the identity
$\tilde{\rho}=\frac{1}{\lambda}\sum_{r}\,\Pi_{r}$ (4.8)
[…]
When $\rho$ is defined through the probability distribution $\bar{p}_{r}=d_{r}^{2}/|G|\,,\,\forall\,r$, the entropic order parameter vanishes. In this case, the probability distribution $\bar{p}_{r}$ defines a symmetric phase [34]. In this same line, one might consider perturbations around this symmetric phase as density matrices $\rho$ given in terms of
$p_{r}$ $=\bar{p}_{r}+\delta p_{r}\,\quad{\rm s.t}\,\quad\sum_{r}\,\delta p_{r}=0\,$, (4.9)
where it is important to note that these patterns may occur due to the presence of non-vanishing one-point functions of non-local global symmetry operators, the only operators that take different expectation values in different phases [34, 58].
It is easy to find that for the case $\delta p_{r}\ll 1$, the entropic order parameter is given by
$S_{\mathcal{A}}(\rho|\tilde{\rho})\sim\sum_{r}\,(\delta p_{r})^{2}\,$,
in accordance with the positivity of the relative entropy. This is also true for the case of non-invertible symmetries that will be treated below.
Said this, it is expected that the behavior of the entropic order parameter behaves quite differently for scenarios where $\delta\,p_{r}\sim\bar{p}_{r}$. For example, in the case of the group $\mathbb{Z}_{2}$ we have
$\{{p\}}$ $=\{1/2,1/2\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=0\,$
$\{{p\}}$ $=\{1,0\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 2\,$
$\{{p\}}$ $=\{0,1\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 2\,$
4.2 The toric code.
In the toric code, the set of anyons is given by $\mathcal{A}=\{1,Y,Z,X\}$ with quantum dimensions $d_{1}=d_{Y}=d_{Z}=d_{X}=1$. The nontrivial fusion rules are those of Rep($\mathbb{Z}_{2}$): $X\otimes X=Y\otimes Y=Z\otimes Z=1,X\otimes Y=Z,X\otimes Z=Y,Y\otimes Z=X$. In standard physics notation the anyons are labeled as $Y=e$ (electric charge at vertex), $Z=m$ (magnetic flux through a plaquette) and $X=\psi$ (fermionic bound state of $e$ and $m$).
We consider the pattern of symmetry breaking induced by the algebra $\mathbb{A}=1+Y$. This leads to a condensed theory $\mathcal{T}=\{\phi\,,\mathbb{X}\}\cong{\rm Vec}(\mathbb{Z}_{2})$ with quantum dimensions $d_{\phi}=d_{\mathbb{X}}=1$. The figure, represents the principal graph of the subfactor associated with gauging the condensable algebra $1+Y$ in the fusion category of the toric code  [^4] The principal graph is a tool used in subfactor theory. The principal graph encodes the fusion rules of irreducible $\mathbb{A}$-modules arising from the inclusion of subfactors associated with anyon condensation. It is a bipartite graph where vertices represent bimodules, and edges represent fusion multiplicities, that is, what we call branching coefficients. It is a bipartite graph that, for finite-depth subfactors, stabilizes to a finite graph. .
[…]
Figure 1: Principal graph for the symmetry breaking pattern induced by the algebra $\mathbb{A}=1+Y$ in the toric code $\mathcal{A}\cong$ Rep($\mathbb{Z}_{2}$). This leads to a condensed theory $\mathcal{T}\cong{\rm Vec}(\mathbb{Z}_{2})$
From the figure, we observe that the only nontrivial branching coefficients are $n^{\phi}_{1}=n^{\phi}_{Y}=1$ and $n^{\mathbb{X}}_{X}=n^{\mathbb{X}}_{Z}=1$. The index of the subfactor inclusion is then given by
$\lambda=\sum_{a\in\mathcal{A}}\,n^{\phi}_{a}\,d_{a}=n^{\phi}_{1}d_{1}+n^{\phi}_{Y}\,d_{Y}=2\,.$ (4.10)
With this, we can derive the matrix $M^{\prime}$ to obtain
[…]
$=\log\lambda-H(p)-(p_{1}+p_{Y})\log(p_{1}+p_{Y})-(p_{X}+p_{Z})\log(p_{X}+p_{Z})\,.$
Some examples of initial states $\rho\in\mathcal{A}$ with their corresponding probabilities $\{{p\}}$ and the value of the entropic order parameter that measures the degree of information loss after condensation are,
$\{{p\}}$ $=\{1/2,1/2,0,0\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=0\,$
$\{{p\}}$ $=\{1/2,0,0,1/2\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 2\,$
$\{{p\}}$ $=\{1/3,1/3,1/3,0\}+{\rm perms}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\frac{1}{3}\log 2\,.$
4.3 Category Rep(S3).
For this category $\mathcal{A}=\{1,X,Y\}$ with quantum dimensions $\{d_{1}=1,d_{X}=1,d_{Y}=2\}$. The non trivial fusion rules are: $X\otimes X=1,X\otimes Y=Y,Y\otimes Y=1\oplus X\oplus Y$.
  1. 1.
    Condensable Algebra $\mathbb{A}=1+X$
    This condensable algebra generates a new vacuum in which the only nontrivial anyonic excitation that is condensed is $X$. Thus, the pattern of symmetry breaking induced by $\mathbb{A}$ implies $\mathcal{T}=\{\phi,t_{1},t_{2}\}\cong{\rm Vec}(\mathbb{Z}_{3})$ with $\{d_{\phi}=1,d_{t_{1}}=1,d_{t_{2}}=1\}$
[…]
    Figure 2: Principal graph for the symmetry breaking pattern induced by the algebra $\mathbb{A}=1+X$ in $\mathcal{A}\cong$ Rep(S3). This leads to a condensed theory $\mathcal{T}\cong{\rm Vec}(\mathbb{Z}_{3})$
    In this case the only nontrivial branching coefficients are $n^{\phi}_{1}=n^{\phi}_{X}=1$ and $n^{t_{1}}_{Y}=n^{t_{2}}_{Y}=1$ and the index of the inclusion is
    $\lambda=\sum_{a\in\mathcal{A}}\,n^{\phi}_{a}\,d_{a}=n^{\phi}_{1}d_{1}+n^{\phi}_{X}\,d_{X}=2\,.$ (4.12)
    With this
    $S_{\mathcal{A}}(\rho|\,\tilde{\rho})=\log\lambda-H(p)-\sum_{r\in\{1,X\}}p_{r}\,\log(p_{1}+p_{X})-p_{Y}\,\log{2\ p_{Y}}$ (4.13)
    Examples of initial states $\rho\in\mathcal{A}$ and the value of the entropic order parameter are
    $\{{p\}}$ $=\{1,0,0\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 2\,$
    $\{{p\}}$ $=\{1/2,1/2,0\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=0\,$
    $\{{p\}}$ $=\{1/2,0,1/2\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\frac{1}{2}\log 2\,$
    $\{{p\}}$ $=\{0,1/2,1/2\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\frac{1}{2}\log 2\,$
    $\{{p\}}$ $=\{1/3,1/3,1/3\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=0.$
  2. 2.
    Condensable Algebra $\mathbb{A}=1+Y$
    This condensable algebra generates a new vacuum in which the only nontrivial anyonic excitation that is condensed is $Y$. The pattern of symmetry breaking induced by $\mathbb{A}$ implies $\mathcal{T}=\{\phi,\mathbb{X}\}\cong{\rm Vec}(\mathbb{Z}_{2})$ with $\{d_{\phi}=1,d_{\mathbb{X}}=1\}$
[…]
    Figure 3: Principal graph for the symmetry breaking pattern induced by the algebra $\mathbb{A}=1+Y$ in $\mathcal{A}\cong$ Rep(S3). This leads to a condensed theory $\mathcal{T}\cong{\rm Vec}(\mathbb{Z}_{2})$
    In this case the only nontrivial branching coefficients are $n^{\phi}_{1}=n^{\phi}_{Y}=1$ and $n^{\mathbb{X}}_{X}=n^{\mathbb{X}}_{Y}=1$ and the index of the inclusion is
    $\lambda=\sum_{a\in\mathcal{A}}\,n^{\phi}_{a}\,d_{a}=n^{\phi}_{1}d_{1}+n^{\phi}_{Y}\,d_{Y}=3\,.$ (4.14)
    With this
    $S_{\mathcal{A}}(\rho|\tilde{\rho})=\log\lambda-H(p)-p_{1}\,\log\left(p_{1}+\frac{p_{Y}}{2}\right)-p_{X}\,\log\left(p_{X}+\frac{p_{Y}}{2}\right)-p_{Y}\,\log 2,$ (4.15)
    Examples of initial states $\rho\in\mathcal{A}$ and the value of the entropic order parameter are
    $\{{p\}}$ $=\{1,0,0\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 3\,$
    $\{{p\}}$ $=\{1/2,1/2,0\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 3\,$
    $\{{p\}}$ $=\{1/2,0,1/2\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\frac{1}{2}\log 3/2\,$
    $\{{p\}}$ $=\{0,1/2,1/2\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\frac{1}{2}\log 3/2\,$
    $\{{p\}}$ $=\{1/3,1/3,1/3\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\frac{1}{3}\log 2.$
  3. 3.
    Condensable Algebra $\mathbb{A}=1+X+2Y$
    This is a Lagrangian algebra. As in the case of the group, the pattern of symmetry breaking induced by $\mathbb{A}$ yields a new vacuum $\phi$ in which all the nontrivial anyonic charges $a$ have condensed, and, this implies that $\mathcal{T}=\{\phi\}\cong{\rm Vec}(\mathbb{C})$ with $\{d_{\phi}=1\}$
[…]
    Figure 4: Principal graph for the symmetry breaking pattern induced by the algebra $\mathbb{A}=1+X+2Y$ in $\mathcal{A}\cong$ Rep(S3). This leads to a trivial condensed theory $\mathcal{T}\cong{\rm Vec}(\mathbb{C})$
    In this case the only nontrivial branching coefficients are $n^{\phi}_{1}=n^{\phi}_{X}=1$ and $n^{\phi}_{Y}=2$ and the index of the inclusion is
    $\lambda=\sum_{a\in\mathcal{A}}\,n^{\phi}_{a}\,d_{a}=n^{\phi}_{1}d_{1}+n^{\phi}_{X}\,d_{X}+n^{\phi}_{Y}\,d_{Y}=6\,.$ (4.16)
    With this
    $S_{\mathcal{A}}(\rho|\tilde{\rho})=\log\lambda-H(p)-2\,p_{Y}\,\log 2\,$ (4.17)
    Some examples of initial states $\rho\in\mathcal{A}$ and the value of the entropic order parameter are
    $\{{p\}}$ $=\{1,0,0\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 6\,$
    $\{{p\}}$ $=\{1/2,1/2,0\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 3\,$
    $\{{p\}}$ $=\{1/2,0,1/2\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 3/2\,$
    $\{{p\}}$ $=\{0,1/2,1/2\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\log 3/2\,$
    $\{{p\}}$ $=\{1/3,1/3,1/3\}\implies S_{\mathcal{A}}(\{p\}|\{\tilde{p}\})=\frac{1}{3}\log 2.$
5 Conclusions
In this work, we have introduced an operator–algebraic and information–theoretic framework to study symmetry-breaking patterns of generalized symmetries in two spatial dimensions. The central idea is to model the restriction and lifting operations associated with a condensable Frobenius algebra by means of conditional expectations and coarse–graining maps. These maps admit explicit Kraus representations determined by the branching coefficients of the condensation, making the construction transparent and computationally tractable.
From this setting, we extracted an information–theoretic order parameter: the quantum relative entropy between a state and its lifted image. This quantity measures the loss of distinguishability induced by condensation, and is bounded universally by the logarithm of the Jones index, which coincides with the quantum dimension of the condensate. In this way, the amount of information erased in the restriction process is directly tied to a fundamental invariant of the subfactor. This gives a precise bridge between algebraic structures underlying anyon condensation and entropic measures familiar from quantum information theory.
[…]
Our results open several directions for future work. On the mathematical side, it would be interesting to extend the present finite–dimensional analysis to type II/III settings, where modular theory plays a central role. On the physics side, the relative entropy order parameter could be tested in lattice realizations of topological phases, or connected to symmetry topological field theory (SymTFT) invariants. Finally, the interplay between condensation, anomalies, and entropic quantities deserves further exploration.
In summary, the combination of subfactor theory, Frobenius algebras, and quantum information tools provides a powerful and versatile language for understanding generalized symmetry breaking and anyon condensation, yielding both conceptual insights and concrete computational tools.
Acknowledgements
We thank Marina Huerta, Horacio Casini, and Javier M. Magán for the stimulating discussions and their kind invitation to participate in the workshop “Anomalies, Topology and Quantum Information in Field Theory and Condensed Matter Physics” (ICTP-SAIFR, São Paulo, Brazil, June 16–27, 2025). J.M.-V. thanks the financial support of Spanish MINECO grant PID2024-155685NB-C22 and Fundación Séneca de la Región de Murcia FSRM/10.13039/100007801(22581/PI/24). G.S. and H-C.Z. acknowledge financial support from the Spanish MINECO grant PID2021-127726NB-I00, the CSIC Research Platform on Quantum Technologies PTI-001 and the QUANTUM ENIA project Quantum Spain funded through the RTRP-Next Generation program under the framework of the Digital Spain 2026 Agenda.
[…]
Appendix A Idempotence and Bimodularity of $\varepsilon_{\mathbb{A}}$
In this appendix we prove two central properties of the conditional expectation $\varepsilon_{\mathbb{A}}$ defining the inclusion $\mathcal{T}\subset\mathcal{A}$: the idempotence, which amounts to say that $\varepsilon_{\mathbb{A}}$ is a projector from $\mathcal{A}$ to $\mathcal{T}$; and the bimodule property.
Idempotence.
We prove the idempotence of $\varepsilon_{\mathbb{A}}$ by computing $\varepsilon_{\mathbb{A}}\left(\tilde{\rho}\right)\in\mathcal{T}$ with $\tilde{\rho}=\alpha(\rho_{\mathbb{A}})$. Graphically this amounts to
$\rho$$\rho_{\mathbb{A}}$$\tilde{\rho}$${\varepsilon_{\mathbb{A}}\left(\tilde{\rho}\right)}$$\scriptstyle{\varepsilon_{\mathbb{A}}}$$\scriptstyle{\alpha}$$\scriptstyle{\varepsilon_{\mathbb{A}}}$
[…]
$\varepsilon_{\mathbb{A}}(\rho)$ $=\sum_{t\in\mathcal{T}}\,p_{t}\,\Pi_{t}\,,\quad p_{t}=\sum_{a\in\mathcal{A}}\,n^{t}_{a}\left(\frac{d_{t}}{d_{a}}\right)\,p_{a}\,$, (A.1)
and
$\tilde{\rho}$ $=\sum_{a\in\mathcal{A}}\,\widetilde{p}_{a}\,\Pi_{a}\,,\quad\widetilde{p}_{a}=\frac{1}{\lambda}\sum_{t\in\mathcal{T}}\,n^{t}_{a}\left(\frac{d_{a}}{d_{t}}\right)\,p_{t}\,.$ (A.2)
Therefore, we find that
[…]
The equality implying idempotence holds if $p_{t}=\tilde{p}_{t},\,\forall\,t\in\mathcal{T}$, which immediately follows from Eqs. (3.8a) and (3.8b).
Bimodule property.
We recall that this refers to
$\varepsilon_{\mathbb{A}}({\tilde{p}}\,{m}\,{\tilde{q}})={p}\,\varepsilon_{\mathbb{A}}({m})\,{q}\,\quad\forall\,{p},{q}\in\mathcal{T},{m}\in\mathcal{A},$
where the tildes refer to the lifted versions of operators in $\mathcal{T}$. To prove this property, we first focus on the first term where
${\tilde{p}}$ $=\sum_{b\in\mathcal{A}}\,{\tilde{p}}_{b}\,\Pi_{b}\,,\quad{\tilde{q}}=\sum_{c\in\mathcal{A}}\,{\tilde{q}}_{c}\,\Pi_{c}\,,\quad{m}=\sum_{d\in\mathcal{A}}\,{m}_{d}\,\Pi_{d}\,.$ (A.4)
Now, we note that for $a\in\mathcal{A}$,
$\mathbb{K}_{a}\,{\tilde{p}}$ $=\sum_{t\in\mathcal{T}}\,\sqrt{\lambda^{t}_{a}}\,\Lambda_{t,a}\left(\sum_{b\in\mathcal{A}}\,{\tilde{p}}_{b}\,\Pi_{b}\right)=\sum_{t\in\mathcal{T}}\,\sum_{b\in\mathcal{A}}\,\sqrt{\lambda^{t}_{a}}\,{\tilde{p}}_{b}\,\Lambda_{t,a}\,\Pi_{b}$ (A.5)
$=\sum_{t\in\mathcal{T}}\,\sqrt{\lambda^{t}_{a}}\,{\tilde{p}}_{a}\,\Lambda_{t,a}\,$,
where $\lambda^{t}_{a}=n^{t}_{a}\,\left(\frac{d_{t}}{d_{a}}\right)$. Similarly,
${\tilde{q}}\,\mathbb{K}^{\dagger}_{a}=\sum_{t\in\mathcal{T}}\,\sqrt{\lambda^{t}_{a}}\,{\tilde{q}}_{a}\,\Lambda^{\dagger}_{t,a}\,.$ (A.6)
With this,
[…]
$=\sum_{t\in\mathcal{T}}\,\left(\sum_{a\in\mathcal{A}}\,\lambda^{t}_{a}\,\left({\tilde{p}}_{a}{\tilde{q}}_{a}\,{m}_{a}\right)\right)\,\Pi_{t}=\sum_{t\in\mathcal{T}}\,c_{t}\,\Pi_{t}\,$,
with
$c_{t}=\sum_{a\in\mathcal{A}}\,\lambda^{t}_{a}\,\left({\tilde{p}}_{a}{\tilde{q}}_{a}\,{m}_{a}\right)\,.$ (A.8)
Now we focus on the second term where
${p}$ $=\sum_{r\in\mathcal{T}}\,{p}_{r}\,\Pi_{r}\,,\quad{q}=\sum_{s\in\mathcal{T}}\,{q}_{s}\,\Pi_{s}\,,\quad\varepsilon_{\mathbb{A}}({m})=\sum_{t\in\mathcal{T}}\,{m}_{t}\,\Pi_{t}\,$, (A.9)
in such a way that
${p}\,\varepsilon_{\mathbb{A}}({m})\,{q}$ $=\sum_{t,r,s\in\mathcal{T}}\,\left({p}_{r}\,{q}_{s}\,{m}_{t}\right)\,\delta_{rs}\delta_{st}\,\Pi_{t}$ (A.10)
$=\sum_{t\in\mathcal{T}}\,\left({p}_{t}\,{q}_{t}\,{m}_{t}\right)\,\Pi_{t}=\sum_{t\in\mathcal{T}}\,c_{t}\,\Pi_{t}\,$,
where the last equality uses Eqs. (3.18) and (A.8), thus proving the bimodule property of $\varepsilon_{\mathbb{A}}$.
Appendix B Dualities and equivalent condensations
Morita equivalence between condensable algebras determines whether two distinct algebraic structures (describing anyon condensation processes) lead to the same topological phase after condensation [59]. Physically, this equivalence captures when different "vacuum redefinitions" via condensation produce identical long-range topological behavior.
Concretely, two condensable algebras $\mathbb{A}$ and $\mathbb{B}$ are Morita equivalent if their categories of local modules-the resulting topological phases-represented by $\mathcal{U}$ in our language, are equivalent as modular tensor categories. While here we are only focusing on the step $\mathcal{A}\to\mathcal{T}$ we note that the equivalence relationship (’Morita equivalence’) between different anyon condensation processes turns out to be just, or is mediated by a duality in the original theory.
We describe here how the existence of a duality in the original theory implies the equivalence of two different anyon condensation processes described by two different condensable algebras $\mathbb{A}$ and $\mathbb{B}$ having in mind that Condensable algebra $\Longleftrightarrow$ Restriction/Lifting Map
First, we consider a duality map in the original theory that may simply be expressed in terms of a permuting symmetry between some anyons $a\in\mathcal{A}$. An example of this is the electric-magnetic duality in the toric code that we will comment later.
Then, if exists, let us assume that the permuting symmetry is mediated by the operator
$\mathrm{Q}=\sum_{a,b\in\mathcal{A}}\,q_{ab}\,X_{ab}\,$, (B.1)
where operators $X_{ab}$ are defined through
$X_{ab}\,\Pi_{b}\,X_{ab}^{\dagger}=\Pi_{a}\,.$ (B.2)
In this setting, we will also consider two different anyon condensation processes that are defined through the branching coefficients,
$(n^{t}_{a})_{\mathbb{A}}$ $\Longleftrightarrow{\rm Algebra}\,\mathbb{A}$
$(n^{t}_{a})_{\mathbb{B}}$ $\Longleftrightarrow{\rm Algebra}\,\mathbb{B}$
and the corresponding conditional expectations
$\mathbb{K}^{\mathbb{A}}_{a}$ $=\sum_{t\in\mathcal{T}}\,\lambda_{ta}^{\mathbb{A}}\,\Lambda_{ta}\,,\quad\lambda_{ta}^{\mathbb{A}}=\sqrt{(n^{t}_{a})_{\mathbb{A}}\left(\frac{d_{t}}{d_{a}}\right)}\,$ (B.3)
$\mathbb{K}^{\mathbb{B}}_{a}$ $=\sum_{t\in\mathcal{T}}\,\lambda_{ta}^{\mathbb{B}}\,\Lambda_{ta}\,,\quad\lambda_{ta}^{\mathbb{B}}=\sqrt{(n^{t}_{a})_{\mathbb{B}}\left(\frac{d_{t}}{d_{a}}\right)}\,.$
Then, it is possible to show that
$\varepsilon_{\mathbb{A}}(\mathrm{Q}\,\rho\,\mathrm{Q}^{\dagger})\equiv\varepsilon_{\mathbb{B}}(\rho)$ (B.4)
and
$\lambda_{ta}^{\mathbb{B}}=\sum_{b\in\mathcal{A}}\,q_{ab}\,\lambda_{tb}^{\mathbb{A}}$ (B.5)
In order to illustrate the idea, let us consider the toric code. As commented in the main text, we first consider the pattern of symmetry breaking induced by the algebra $\mathbb{A}=1+Y$ which implies that $\mathcal{T}=\{\phi\,,\mathbb{X}\}\cong{\rm Vec}(\mathbb{Z}_{2})$ with $\{d_{\phi}=d_{\mathbb{X}}=1\}$.
For convenience, we show again the the principal graph of the subfactor inclusion induced by $\mathbb{A}$
[…]
Figure 5: Principal graph for the symmetry breaking pattern induced by the algebra $\mathbb{A}=1+Y$ in the toric code $\mathcal{A}\cong$ Rep($\mathbb{Z}_{2}$). This leads to a condensed theory $\mathcal{T}\cong{\rm Vec}(\mathbb{Z}_{2})$
The only non-trivial branching coefficients are $n^{\phi}_{1}=n^{\phi}_{Y}=1$ and $n^{\mathbb{X}}_{X}=n^{\mathbb{X}}_{Z}=1$. This refers to
$\lambda_{\phi 1}^{\mathbb{A}}=\lambda_{\phi Y}^{\mathbb{A}}=1=\lambda_{\mathbb{X}X}^{\mathbb{A}}=\lambda_{\mathbb{X}Z}^{\mathbb{A}}=1\,$, (B.6)
which implies
$\varepsilon_{\mathbb{A}}(\rho)$ $=p_{\phi}\,\Pi_{\phi}+p_{\mathbb{X}}\,\Pi_{\mathbb{X}}$ (B.7)
$p_{\phi}$ $=p_{1}+p_{Y}\,$
$p_{\mathbb{X}}$ $=p_{X}+p_{Z}$
On the other hand, we consider the condensable Algebra $\mathbb{B}=1+Z$ whose pattern of symmetry breaking gives $\mathcal{T}=\{\phi\,,\mathbb{X}\}\cong{\rm Vec}(\mathbb{Z}_{2})$ with $\{d_{\phi}=d_{\mathbb{X}}=1\}$ and a principal graph.
[…]
Figure 6: Principal graph for the symmetry breaking pattern induced by the algebra $\mathbb{A}=1+Z$ in the toric code $\mathcal{A}\cong$ Rep($\mathbb{Z}_{2}$). This leads to a condensed theory $\mathcal{T}\cong{\rm Vec}(\mathbb{Z}_{2})$
The only non-trivial branching coefficients are $n^{\phi}_{1}=n^{\phi}_{Z}=1$ and $n^{\mathbb{X}}_{X}=n^{\mathbb{X}}_{Y}=1$. This refers to
$\lambda_{\phi 1}^{\mathbb{B}}=\lambda_{\phi Z}^{\mathbb{B}}=1=\lambda_{\mathbb{X}X}^{\mathbb{B}}=\lambda_{\mathbb{X}Y}^{\mathbb{B}}=1\,$, (B.8)
which implies
$\varepsilon_{\mathbb{B}}(\rho)$ $=p_{\phi}\,\Pi_{\phi}+p_{\mathbb{X}}\,\Pi_{\mathbb{X}}$ (B.9)
$p_{\phi}$ $=p_{1}+p_{Z}\,$
$p_{\mathbb{X}}$ $=p_{X}+p_{Y}$
The permutation symmetry of the toric code interchanges $Y\Longleftrightarrow Z$ through the operator.
${\rm Q}=\left(\begin{array}[]{cc}\mathbb{I}_{2}&0\\ 0&X\end{array}\right)\,$, (B.10)
where $X$ is Pauli matrix
$X=\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right)\,$, (B.11)
from which it is straightforward to check that
$\varepsilon_{\mathbb{A}}(\mathrm{Q}\,\rho\,\mathrm{Q}^{\dagger})\equiv\varepsilon_{\mathbb{B}}(\rho)\,.$ (B.12)